OpenStudy (anonymous):

cos(pi-o)+ sin(pi/2 +0)=0 verifying the identity

OpenStudy (anonymous):

theta?

OpenStudy (anonymous):

\[cos(\pi-\theta)+sin({\pi\over2}+\theta)=0~?\]

OpenStudy (whpalmer4):

Think about what happens to the value of \(\cos \theta\) if you add \(\pi\) to \(\theta\). It flops over to the diagonally opposite spot on the unit circle, because the period of the sin and cos functions is \(2\pi\), so adding \(\pi\) takes you halfway around the circle from where you started. Also, \[\cos \theta = \cos -\theta\]which means the left hand term is simply the cos function shifted in phase by \(\pi\) radians or \(180^\circ\). Now, what does adding \(\pi/2\) to the argument of the sin function do? Does the graph of \(\sin(\dfrac{\pi}{2}+\theta)\) resemble the graph of any other basic trig function?